OpenStudy (anonymous):

A circle is represented by the equation shown below. (x - 5)2 + (y + 9)2 = 144 Which statement is true? a. The circle is centered at (-5, -9) and has a diameter of 12. b. The circle is centered at (5, -9) and has a radius of 12. c. The circle is centered at (5, 9) and has a diameter of 12. d. The circle is centered at (5, 9) and has a radius of 12.

OpenStudy (ranga):

(x - h)^2 + (y - k)^2 = r^2 is the equation of a circle with the center at (h, k) and radius = r. Compare this to (x - 5)^2 + (y + 9)^2 = 12^2 What is h, k and r?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

jello lol

OpenStudy (anonymous):

5, -9, 12 ?

OpenStudy (ranga):

Yes.